### v1

# Two dimensional de-Sitter and deformed CFTs

Abstract
[…]
1 Introduction
In a closed universe, ‘the problem of time’ appeared to be the most challenging conceptual as well as technical aspect to formulate a quantum theory of gravity [2]-[4]. In an open universe like AdS, where the ADM Hamiltonian acquires boundary term, one can define asymptotic observers and observables with an extrinsic notion of boundary time. In AdS/CFT, the UV completed boundary CFT thus has a well-defined intrinsic notion of time. Even though the notion of proper boundary time could be different for the action of large diffeomorphism in asymptotically AdS(specially at two and three dimensions), the boundary ‘external’ observers can keep track of the history with their own ADM clock [5]. However, in closed universe like dS, general covariance of GR dramatically contradicts with the quantum mechanical notion of ‘fixed time’ through Wheeler-Dewitt equation: $\hat{H}|\psi\rangle_{uni}=0$ [6], [7]. Here the vanishing Hamiltonian constraints on physical state of the universe would dictates a ‘static universe’ that directly contradicts with our observed experience. In other way, ‘the problem of time’ indirectly suggests that time should be an emergent phenomenon in a quantum theory of gravity in a closed universe.
One of the most compelling resolutions was proposed by Page and Wootters[8]. According to it, for a closed quantum system subject to a Hamiltonian constraint (such as the Wheeler–DeWitt equation), an internal observer can experience dynamics via a ‘relational time,’ arising from the system’s entanglement with a clock subsystem. More recently, this resolution has been recast as an addition of a gravitating observer in a closed universe with a goal towards defining a background independent formulation of quantum gravity[9]-[12]. Of course, by definition the observer should be emergent in the full UV theory. To be more precise, in low energy limit, one may think of the total Hamiltonian of an observed system to be sum of the Hamiltonian of the subsystem $H_{S}$ and the Hamiltonian of the gravitating observer $H_{O}$ [12]. In special cases [^1] When we neglect the space-time fluctuation of gravitational subsystem in a sufficiently low energy., one may take $H_{S}$ (upto a multiplicative constant) as the modular Hamiltonian of the subsystem with respect to a cyclic, separating state vector $|\psi\rangle_{S}$. In this combined system, one can define $|\psi(t)\rangle_{S}=e^{-iH_{S}t}|\psi(0)\rangle_{S}$, where the projected $|\psi(0)\rangle_{S}\equiv\langle\phi(0)|\psi\rangle$ and $|\phi(t)\rangle_{O}=e^{-iH_{O}t}|\phi(0)\rangle_{O}$. In this way, imposing the constraint would tell us the observer measures modular time defined by $H_{S}$ with respect to $|\psi\rangle_{S}$ on the subsystem $S$.
Lately, there has been considerable work [12]-[24] [^2] This is not exhaustive, see the cited works for more. exploring the idea of including the observer as part of the gravitational system in a closed universe following [11]. The most natural setting to investigate this proposal is a closed de Sitter space-time, where a static observer’s accessible region is confined to a static patch bounded by cosmological (as well as black hole) horizons. Apart from a theoretical playground, dS universe has a more direct contact with observational cosmology and the fate of our living universe. The widely studied inflationary regime of cosmology has been approximately described by a metastable dS phase. On the other hand, the present observation of accelerated expansion also suggests that the universe will go through an asymptotic de-sitter phase at late time. If we imagine an astronomical observer of the current time, it will probably see a locally dS static patch in his own clock after sufficiently long time. Hence understanding the huge amount of ‘mysterious’ Hawking-Gibbons entropy of static patch [25] and consequently providing a workable model of ‘static patch holography’ would be a great deal of interest. The most obvious notion may come from the stretched horizon holography, where in the absence of an asymptotic boundary, the vast amount of entropy could be located near the horizon[26],[27]. More intuitively, if one may imagine an observer moving along the timelike stretched horizon, he could effectively store all the degrees of freedom and measure the same horizon entropy [^3] Note that, an interesting feature of the static patch cosmological horizon is it’s observer dependence. Even though, different observer could see different horizon, the area is always be the same.. Nevertheless, only a clean UV theory of static patch effective theory could lead us to the correct and proper understanding of static patch entropy and holography.
By definition, the UV theory should have a low energy limit that describes (semi)classical theory of gravity [^4] This should be a minimal requirement for a consistent UV theory.. In the closed universe, as we mentioned, the UV theory should not have a definite notion of time. Here, in a modest sense, our goal of this note is to find a sensible theory that only describes an observer in the lower dimensional dS static patch in low energy limit. To be more precise, we are interested to study two dimensional effective theory of dS in specific coordinates and it’s possible UV description. With this motivation, we study S-wave dimensional reduction of asymptotic Schwarzchild dS4 geometry in the near nariai near horizon limit to find an effective 2D description of dilatons in section (2). In contrast to JT gravity, we provide an alternative reduction, in which the effective theory can be recast as free massless dilatons propagating in a static patch background enclosed by cosmological and blackhole horizon with same temperature. We quantize the dilatons using some UV dirichlet boundary condition at the stretched horizon and find a finite entropy which we interpret as out-side entropy $S_{out}$ of observers living inside the static patch with a worldline along the stretched horizon. In section (3), we consider S-wave reduction of pure dS3 in 3D Einstein gravity. Unlike a free theory, this effective theory(from any higher dimensional reduction) can be written as dynamical dilatons inside a past Milne wedge. We also observe the dilaton effective theory has an initial potential dominance and late time kinetic dominance, providing us a simple toy model for inflation in 2D. We relate it to a critical quench problem and then discuss a boundary non-preserving deformed CFT Hamiltonian on a strip in section 2.1. The ultimate fate of the constructed past Milne vacuum is shown to be unstable at (finite but) late time, where the kinetic term dominates over potential term in section (4.2). Since Milne patch is analytically continued to static patch with thermal Euclidean time circle, we immediately constructs a sensible worldsheet description of (extended) static patch [^5] Throughout the note, we refer ‘extended static patch’ as the dimensionally reduced Nariai black hole enclosed by black hole and cosmological horizon at same temperature. in terms of deformed CFT Hamiltonian on a ring in section (4). We further study it’s modular quantization and connect the thermal entropy with entanglement entropy of static patch vacuum in section (5.1) and (5.2). The boundary condition on the stretched horizon is an emergent requirement of this quantization, as we also noticed in AdS [28],[1]. We also discuss our proposed alternative S-wave reduction in 4D magnetically charged RN AdS4 blackhole in the near horizon near extremal limit and find free massless dilatons coupled to boundary Schwarzian as an effective description in appendix (A). This provides a closure to our claim in a previous paper [1] on the UV completion of JT coupled to $c=\mathcal{O}(1)$ CFT as a worldsheet deformed CFT in a more concrete way.
However, in this modest approach, the questions regarding the observer dependent choice of UV time is still obscure, since our description of worldsheet CFT already knows the ‘time’ of what observer will see in a static patch. We know in dS, different patches describe different notion of time and hence finding a global vacuum is inherently a difficult problem [^6] See [10] and the discussion in [21] from algebra of observables point of view.. We end our note with some comments on it in section 5 along with possible future direction.
2 Dimensional Reduction to 2D de-Sitter in near-Nariai limit
Here we review the dimensional reduction of pure Einstein gravity in 4D to two-dimensional JT gravity in the near-Nariai limit following [29]. Consider the following metric ansatz of the class of spherically symmetric asymptotically 4D de Sitter metric:
$ds^{2}_{4}=g_{\mu\nu}dx^{\mu}dx^{\nu}+r_{N}^{2}\phi(x)d\Omega^{2}_{2}$ (2.1)
Here, $g_{\mu\nu}$ is an unspecified two-dimensional metric with $\mu,\nu=(0,1)$ and $\phi(x)$ is the dilaton field. The corresponding four-dimensional Einstein-Hilbert action is
$I_{4}=\frac{1}{16\pi G_{4}}\int_{M_{4}}d^{4}X\sqrt{-\hat{g}_{4}}(\hat{R}_{4}-2\Lambda);\;\Lambda=\frac{3}{L_{4}^{2}}$ (2.2)
Here $X=(x_{\mu},\Omega_{2})$. $\hat{g}_{4}$ and $\hat{R}_{4}$ denote the determinant of the four-dimensional metric (2.1) and the corresponding Ricci scalar, respectively. $\hat{R}_{4}$ also reduces to the following form [29] :
$\hat{R}_{4}=R_{2}+\frac{2}{r_{N}^{2}\phi}+\frac{(\nabla\phi)^{2}}{2\phi^{2}}-\frac{2}{\phi}\nabla^{2}\phi$ (2.3)
Here, $R_{2}$ and $\nabla$ refer to a two-dimensional Ricci scalar and the covariant derivative with respect to the 2D metric $g_{\mu\nu}(x)$. Also, because of spherical reduction we have,
$\int_{M_{4}}d^{4}X\sqrt{-\hat{g}_{4}}=r_{N}^{2}\Omega_{2}\int_{M_{2}}d^{2}x\sqrt{-g}\phi$ (2.4)
By defining, $\frac{1}{G_{2}}\equiv\frac{r_{N}^{2}\Omega_{2}}{G_{4}}$ and by using (2.3) and (2.4), the action (2.2) reduces to the following 2D action
$I_{4}=\frac{1}{16\pi G_{2}}\int_{M_{2}}d^{2}x\sqrt{-g}\left(\phi R_{2}-2\Lambda\phi+\frac{2}{r_{N}^{2}}+\frac{(\nabla\phi)^{2}}{2\phi}-2\nabla^{2}\phi\right)$ (2.5)
In order to obtain a JT gravity action, the kinetic term must be dropped out. This can be achieved by a Weyl rescaling of the metric $g_{\mu\nu}\rightarrow\omega^{2}(x)\bar{g}_{\mu\nu}$. Under such Weyl rescaling, the metric determinant and $R_{2}$ transform as:
$\sqrt{-g}=\omega^{2}\sqrt{-\bar{g}}\;;\;R_{2}=\frac{\bar{R}_{2}}{\omega^{2}}-\frac{2\nabla^{2}\omega}{\omega^{3}}+\frac{2(\nabla\omega)^{2}}{\omega^{4}}.$ (2.6)
Thus putting (2.6) back into (2.5), we end up with
$I_{4}=\frac{1}{16\pi G_{2}}\int_{\bar{M}_{2}}d^{2}x\sqrt{-\bar{g}}\left[\phi\bar{R}_{2}-\frac{2\phi\nabla^{2}\omega}{\omega}+\frac{2\phi(\nabla\omega)^{2}}{\omega^{2}}-2\Lambda\phi\omega^{2}+\frac{2\omega^{2}}{r_{N}^{2}}+\frac{(\nabla\phi)^{2}}{2\phi}-2\nabla^{2}\phi\right]$ (2.7)
Note that, here the covariant derivative $\nabla$ is taken with respect to $\bar{g}_{\mu\nu}$. Now we can write
$\frac{2\phi(\nabla\omega)^{2}}{\omega^{2}}=\nabla^{\mu}(\frac{2\phi}{\omega}\nabla_{\mu}\omega)-\frac{2}{\omega}(\nabla^{\mu}\phi)(\nabla_{\mu}\omega)-\frac{2\phi}{\omega}\nabla^{2}\omega-2\phi\omega\nabla^{\mu}(\frac{1}{\omega^{2}})\nabla_{\mu}\omega$
$\implies\frac{2\phi(\nabla\omega)^{2}}{\omega^{2}}=\frac{2\phi}{\omega}\nabla^{2}\omega+\frac{2}{\omega}(\nabla^{\mu}\phi)(\nabla_{\mu}\omega)-\nabla^{\mu}(\frac{2\phi}{\omega}\nabla_{\mu}\omega)$ (2.8)
Inserting the above identity in (2.7), we have
$I_{4}=\frac{1}{16\pi G_{2}}\int_{\bar{M}_{2}}d^{2}x\sqrt{-\bar{g}}\left[\phi\bar{R}_{2}+\frac{2}{\omega}(\nabla^{\mu}\phi)(\nabla_{\mu}\omega)-2\Lambda\phi\omega^{2}+\frac{2\omega^{2}}{r_{N}^{2}}+\frac{(\nabla\phi)^{2}}{2\phi}-2\nabla^{2}\phi-2\nabla^{\mu}(\frac{\phi}{\omega}\nabla_{\mu}\omega)\right]$ (2.9)
If we now choose $\omega=\frac{\alpha}{\phi}$, we can eliminate the total derivative term as well as the kinetic term, since
$\frac{2}{\omega}(\nabla^{\mu}\phi)(\nabla_{\mu}\omega)|_{\omega=\frac{\alpha}{\phi}}=-\frac{(\nabla\phi)^{2}}{2\phi};\;2\nabla^{\mu}(\frac{\phi}{\omega}\nabla_{\mu}\omega)|_{\omega=\frac{\alpha}{\phi}}=-2\nabla^{2}\phi$ (2.10)
Hence the choice simplifies the effective 2D action as
$I_{4}=\frac{1}{16\pi G_{2}}\int_{\bar{M}_{2}}d^{2}x\sqrt{-\bar{g}}\left[\phi\bar{R}_{2}-\frac{2\Lambda\alpha^{2}}{\phi}+\frac{2\alpha^{2}}{r_{N}^{2}\phi^{2}}\right]$ (2.11)
Note that, this choice of scaling factor is slightly different than [29], in which the total derivative terms in the effective action are getting canceled with the boundary Gibbons-Hawking term. However, here we are not adding a GHY boundary term in a closed universe setting. Hence the potential term will also be different from theirs. However the end result is similar. Using $\Lambda=\frac{3}{L_{4}^{2}}$, the Nairai black hole radius $r_{N}=\frac{L_{4}}{\sqrt{3}}$ and choosing $\alpha=\frac{1}{\sqrt{3}}$ we finally obtain
$I_{4}=\frac{1}{16\pi G_{2}}\int_{\bar{M}_{2}}d^{2}x\sqrt{-\bar{g}}\left[\phi\bar{R}_{2}+U(\phi)\right];\;\text{where}\;U(\phi)=\frac{2}{L_{4}^{2}}\left(-\frac{1}{\phi}+\frac{1}{\phi^{2}}\right)$ (2.12)
Once again,
$U(\phi=1)=0,\;\frac{dU}{d\phi}|_{\phi=1}=-\frac{2}{L^{2}_{4}}$ (2.13)
Thus by expanding $\phi=\phi_{0}+\Phi$ with $\phi_{0}=1$, the leading order term of $I_{4}$ provides the effective JT gravity action as
$I_{4}\approx I_{JT}=\frac{1}{16\pi G_{2}}\int_{\bar{M}_{2}}d^{2}x\sqrt{-\bar{g}}\phi_{0}\bar{R}_{2}+\frac{1}{16\pi G_{2}}\int_{\bar{M}_{2}}d^{2}x\sqrt{-\bar{g}}\Phi\left[\bar{R}_{2}-\frac{2}{L_{4}^{2}}\right]$ (2.14)
One can readily recognize the 1st term as the topological Gauss-Bonnet term without boundary, whereas the second term belongs to the standard JT gravity with positive cosmological constant. One can connect it with the standard reduction of JT gravity from 4D Schwarzchild de-Sitter(SdS) black holes in near Nariai limit. The metric of 4D SdS black hole is the following [29]:
$ds^{4}=-f(r)dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}d\Omega^{2}_{2};\;\text{where}\;f(r)=1-\frac{r^{2}}{L_{4}^{2}}-\frac{16\pi G_{4}M}{2\Omega_{2}r}=\frac{(r-r_{h})(r_{c}-r)(r+r_{h}+r_{c})}{rL_{4}^{2}}.$ (2.15)
$M$ is the black hole mass. Here $f(r)$ has two positive root: $f(r_{h})=f(r_{c})=0$, where $r_{h}$ and $r_{c}$ are black hole and cosmological horizon respectively. Also $r_{h}<r<r_{c}$. In terms of these horizon radius, we can express [29]
$L_{4}^{2}=\frac{r_{c}^{3}-r_{h}^{3}}{r_{c}-r_{h}},\;\frac{16\pi G_{4}M}{2\Omega_{2}}=\frac{r_{h}r_{c}^{3}-r_{h}^{3}r_{c}}{r_{c}^{3}-r_{h}^{3}}.$ (2.16)
[…]
$\tau=\frac{\beta\epsilon t}{2\hat{L}_{4}},\;\rho=\frac{2\hat{L}_{4}}{\beta}\left(\frac{r-r_{h}}{\epsilon}-\frac{\beta}{2}\right),\;\beta=\frac{r_{c}-r_{h}}{\epsilon},\;\hat{L}_{4}=\frac{L_{4}}{\sqrt{3}}=r_{N}.$ (2.17)
[…]
$ds^{2}=-\left(1-\frac{\rho^{2}}{r_{N}^{2}}\right)d\tau^{2}+\left(1-\frac{\rho^{2}}{r_{N}^{2}}\right)^{-1}d\rho^{2}+r_{N}^{2}d\Omega_{2}^{2}$ (2.18)
Here $\rho\in[-r_{N},r_{N}]$. This coincide with (2.1) where $\phi=1$ and $g_{\mu\nu}$ corresponds to 2D de-Sitter in static patch.
We can then study the effective JT action (2.14) and it’s quantization using standard gravitational path integral procedure [30],[31]. The usual approach is to treat the JT gravity itself as an effective theory of 2D gravity and then proceed to the rigorously well-defined gravitational path integral with dilaton boundary condition as an UV input. For instance, one can simply compute the Hartle-Hawking wave function by doing explicit path integral of the form
$\psi_{+}[h]\propto\int_{\bar{g}_{\partial M}=h,\Phi_{\partial M}=\phi_{b}}D\bar{g}_{\mu\nu}D\Phi e^{-I_{JT}}$ (2.19)
This is relevant for studying model of 2D quantum gravity independent of it’s origin in higher dimension. However, this type of path integral is only meaningful over Hartle-Hawking type geometries, where one can specify certain boundary condition on metric as well as dilaton [31] at conformal infinity. However when considering the static patch, one can recover the total entropy as follows [29]-[32]:
$S=\frac{2\phi_{0}}{4G_{2}}+\frac{\Phi_{bh}}{4G_{2}}+\frac{\Phi_{c}}{4G_{2}}$ (2.20)
Since $\Phi_{bh}=-\Phi_{c}$, the net entropy becomes the entropy of Nariai black hole, i.e. $S=S_{N}=\frac{2\phi_{0}}{4G_{2}}$. Without adding such boundary GHY term, the path integral becomes trivial as it fixes $\bar{R}_{2}=\frac{2}{L^{2}_{4}}$ and $\bar{g}_{\mu\nu}$ is fixed to be locally dS2 metric. Hence the sole contribution to the entropy coming only from $\phi_{0}$ or Nariai term and the near Nariai correction does not contribute unlike JT in AdS2.
2.1 Alternative dual reduction in the near-Nariai limit
Another approach is to take care of the dimensional reduction and quantizing the effective theory governed by 4D gravitational path integral. Since we use a Weyl rescaling $g_{\mu\nu}=\omega^{2}\hat{g}_{\mu\nu}$ and then choose $\omega$ to be some function of $\Phi$, the effective theory has only dynamical variable $\Phi$ once we fix the background $\hat{g}_{\mu\nu}$ to be that of (2.18) in the near Nariai limit [33]. By exploiting conformal flatness of any two-dimensional metric, we can always make this choice which further restricts the form of $R_{2}$ in terms of dilatons. Since the dilatons are arbitrary at this moment, we can always do such fixing. Thus we can fix $\hat{R}_{2}=\frac{2}{\hat{L}_{4}^{2}}$. Finally we are left with only
$Z=e^{-I_{0}}\int D\Phi e^{-\frac{1}{4\pi G_{2}L^{2}_{4}}\int d\rho d\tau_{E}\Phi(\rho,\tau_{E})}=Ae^{-I_{0}},$
where $I_{0}=\frac{\phi_{0}}{16\pi G_{2}}\int^{2\pi\hat{L}_{4}}_{0}d\tau_{E}\int^{\hat{L}_{4}}_{-\hat{L}_{4}}d\rho\frac{2}{\hat{L}^{2}_{4}}=\frac{\phi_{0}}{2G_{2}},\;\text{and}\;A=\int D\Phi e^{-\frac{1}{4\pi G_{2}L^{2}_{4}}\int d\rho d\tau_{E}\Phi(\rho,\tau_{E})}$ (2.21)
Hence, the entropy is $S=S_{N}+\log(A)$. This suggests, the total entropy may have some correction which is not observed from previous way. However, this is not a useful way to study dynamics of the effective theory. This motivates us to consider an inequivalent or alternative dimensional reduction which may capture another useful and tractable description of near horizon physics in closed universe.
To derive JT action, we have used a particular scaling form $\omega$. In stead, we could choose any other $\omega=f(\phi)$, which may cancel the $\frac{(\nabla\phi)^{2}}{2\phi}$ term in the action (2.9). For instance, choosing $\omega=c\phi^{-\frac{1}{4}}e^{\beta\phi}$ in (2.9) yields
$I^{\prime}_{4}=\frac{1}{16\pi G_{2}}\int_{\bar{M}_{2}}d^{2}x\sqrt{-\bar{g}}\left[\phi\bar{R}_{2}+2\beta(\nabla\phi)^{2}+2\Lambda c^{2}e^{2\beta\phi}\left(\frac{1}{\sqrt{\phi}}-\sqrt{\phi}\right)-\nabla^{2}\phi-2\beta\nabla^{\mu}(\phi\nabla_{\mu}\phi)\right]$ (2.22)
Here $\beta$ is a dimensionful constant. This is a Liouville-type action with an effective potential
$U^{\prime}(\phi)=2\Lambda c^{2}e^{2\beta\phi}\left(\frac{1}{\sqrt{\phi}}-\sqrt{\phi}\right),\;\text{such that:}\;U^{\prime}(\phi=1)=0,\frac{dU^{\prime}}{d\phi}|_{\phi=1}=-2\Lambda c^{2}e^{2\beta}$ (2.23)
If we further choose $\beta=\frac{1}{4}$ and $c=e^{-\frac{1}{4}}$, after expanding $\phi=\phi_{0}+\Phi$ with $\phi_{0}=1$, we end up with
$I^{\prime}_{4}=\frac{\phi_{0}}{16\pi G_{2}}\int_{\bar{M}_{2}}d^{2}x\sqrt{-\bar{g}}\bar{R}_{2}+\frac{1}{16\pi G_{2}}\int_{\bar{M}_{2}}d^{2}x\sqrt{-\bar{g}}\left[\frac{1}{2}(\nabla\Phi)^{2}+\Phi\left(\bar{R}_{2}-\frac{2}{\hat{L}^{2}_{4}}\right)-\frac{3}{2}\nabla^{2}\Phi-\frac{1}{2}\nabla^{\mu}(\Phi\nabla_{\mu}\Phi)\right]$ (2.24)
Using the same logic which we explained earlier, we could always fix the two-dimensional $\bar{g}_{\mu\nu}$ to be the static patch metric with $\bar{R}_{2}=\frac{2}{\hat{L}^{2}_{4}}$. Note that by doing this one would find a relation between $R_{2},\frac{2}{\hat{L}^{2}}$ and $\phi$. Hence, by varying $\phi$, we could in principle get all possible $g_{\mu\nu}$ to be consistent with (2.1). Thus, we end up with
$I^{\prime}_{4}=I_{0}+\frac{1}{16\pi G_{2}}\int d\rho d\tau\left[\frac{1}{2}(\nabla\Phi)^{2}-\frac{3}{2}\nabla^{2}\Phi-\frac{1}{2}\nabla^{\mu}(\Phi\nabla_{\mu}\Phi)\right]$ (2.25)
This is a free scalar field action upto two total derivative terms. Those terms can be vanishing if we use the following boundary condition at the two horizons and the usual periodic boundary condition at $(\tau_{E}=0,2\pi\hat{L}_{4})$ in the Euclidean setting:
$\Phi(\rho,\tau_{E})|_{\rho=\pm\hat{L}_{4}}=0,\;\nabla_{\tau}\Phi(\rho,\tau_{E})|_{\tau_{E}=0}=\nabla_{\tau}\Phi(\rho,\tau_{E})|_{\tau=2\pi\hat{L}_{4}},\;\Phi(\rho,0)=\Phi(\rho,2\pi\hat{L}_{4})$ (2.26)
The periodic boundary condition is natural from the periodic nature of thermal Euclidean time circle. However, the Dirichlet boundary condition on $\Phi$ is non-standard and is not a direct requirement of the classical action. At this classical level, it seems to be ad-hoc to use such boundary condition on dilatons at the horizons(or more correctly on the stretched horizons). However, as we will argue later, one must impose such boundary condition in order to quantize a deformed CFT on a ring which further describes CFT on extended static patch background. Hence we may think of this as an emergent boundary condition originated from certain UV theory having a worldsheet description in terms of deformed CFT on a ring. In this alternative dimensional reduction, the near Nariai dynamics can be recast as free massless scalar field in dS2 static patch background with Dirichlet boundary condition at the horizons. In this reduction, the total entropy is
$S=\frac{\phi_{0}}{2G_{2}}+S_{out}^{\Phi},$ (2.27)
where $S_{out}^{\Phi}$ is the entropy of the fields $\Phi$. We may think of this as a generalized entropy of the extended static patch [^7] In contrast to pure de-Sitter static patch, this contains both black hole and cosmological horizon as reduced from the Nariai limit of SdS black holes in 4D. in dS2, which is well-defined in the algebraic formulation of gravity by including observers in the static patch [9]. The role of the observer should be evident while computing the explicit entropy of dilatons as we do in next section.
2.2 Partition function of massless free scalar fields in a static patch background with stretched horizon
It would be useful to perform a coordinate change to map the static patch into conformally flat spacetime coordinate. Using Tortoise-like coordinate $\theta$ defines as
$d\theta=\frac{d\rho}{1-\frac{\rho^{2}}{r_{N}^{2}}}\implies\left[\frac{\theta}{r_{N}}=\frac{1}{2}\ln\frac{1+\frac{\rho}{r_{N}}}{1-\frac{\rho}{r_{N}}}=\tanh^{-1}\frac{\rho}{r_{N}}\right],$ (2.28)
the metric of the static patch becomes
$ds^{2}=\frac{-d\tau^{2}+d\theta^{2}}{\cosh^{2}(\frac{\theta}{r_{N}})}$ (2.29)
Assuming a separation of variable by exploiting time translational symmetry $\Phi(\theta,\tau)=\psi_{\omega}(\theta)e^{-i\omega\tau}$, the solution of massless Klein-Gordon equation $\nabla^{2}\Phi(\theta,\tau)=0$ is
$\psi_{\omega}(\theta)=Ae^{i\omega\theta}+Be^{-i\omega\theta}$ (2.30)
To impose Dirichlet boundary condition at the horizons, we use brick wall regularization pioneered by t’hooft [34]. We consider streteched horizons located at $\rho=r_{N}-\epsilon$ and $\rho=-r_{N}+\epsilon$. Correspondingly in conformal frame,
$\frac{\theta}{r_{N}}|_{\rho=r_{N}-\epsilon}=\frac{1}{2}\ln(\frac{2r_{N}}{\epsilon})\equiv\Lambda_{c},\;\frac{\theta}{r_{N}}|_{\rho=-r_{N}+\epsilon}=\frac{1}{2}\ln(\frac{\epsilon}{2r_{N}})\equiv-\Lambda_{c}$ (2.31)
Hence imposing Dirichlet condition at stretched horizons, we end up with $A=\pm B$. Choosing A=BA=-B, the solution becomes $\Phi=A\sin(\omega\theta)e^{-i\omega\tau}$. The Dirichlet boundary condition further implies [35]
$\sin\omega r_{N}\Lambda_{c}=0\rightarrow\;\omega_{n}=\frac{2n\pi}{r_{N}\ln(\frac{2r_{N}}{\epsilon})},n\in Z$ (2.32)
This form exactly coincides with the emergent discrete Virasoro modes of UV deformed CFT as we will describe in section 5. Choosing A=BA=-B would give $(2n+1)\pi/2\Lambda_{c}$ which does not coincide with the UV theory requirement. Note that, one could also give same form of normal mode by imposing Neumann boundary condition. Also one can check the boundary term $\partial_{\theta}\Phi(\theta=\Lambda_{c})-\partial_{\theta}\Phi(\theta=-\Lambda_{c})]$ exactly vanishes  [^8] The other term will also be vanishing by Dirichlet condition itself..
To fix the normalization $A$, one can further use Klein-Gordon inner product as in [36] which provides an orthonormal basis of mode functions. This is also needed to construct the Hilbert space for the fields. One can easily check that $A=\frac{1}{\sqrt{2\omega_{n}}}$. To compute the partition function, this normalization is crucial. Since we know the exact normal modes, we can perform now an exact computation of partition function as in [36],[37]  [^9] See also [38].after quantizing the fields. We can write the operator $\Phi$ in terms of normalized modes $\psi_{\omega_{n}}$ and the creation and annihilation operators as:
$\Phi(\theta,\tau)=\sum_{n=0}^{\infty}\frac{1}{\sqrt{2\omega_{n}}}\left[a_{n}\psi_{\omega_{n}}(\theta)e^{-i\omega_{n}\tau}+a_{n}^{\dagger}\psi_{\omega_{n}}(\theta)e^{i\omega_{n}\tau}\right]$ (2.33)
Here $a_{n}$ and $a_{n}^{\dagger}$ satisfies canonical commutation relation and also defines the ‘Boulware vacuum’ $|0\rangle_{B}$ by [39],[36]
$a_{n}|0\rangle_{B}=0$ (2.34)
Hence the normal ordered Hamiltonian is also given by
$:H:=\sum_{n}\omega_{n}a_{n}^{\dagger}a_{n}$ (2.35)
The Hilbert space can now be obtained by considering all eigenstates of the Hamiltonians which os of the known form:
$|n\rangle=\prod_{n}\frac{1}{\sqrt{N_{n}}}(a^{\dagger})^{N_{n}}|0\rangle_{B},\;\text{such that}\;:H:|n\rangle=\sum_{n}(\omega_{n}N_{n})|n\rangle$ (2.36)
[…]
$Z_{N}(\beta_{N})=Tr(e^{-\beta_{N}:H:})=\prod_{n}\sum_{N_{n}}e^{-\beta_{N}\omega_{n}N_{n}}=\prod_{n}\frac{1}{1-e^{-\beta_{N}\omega_{n}}}$ (2.37)
Thus taking $\log$ we get
$\log(Z_{N})=-\sum_{n}\log(1-e^{-\beta_{N}\omega_{n}})$ (2.38)
In $\epsilon\rightarrow 0$ limit, the discrete normal mode spectrum becomes continuous $\omega_{n}\rightarrow\omega$ and we can convert the sum into integral as follows:
$\log(Z_{N})=-\frac{r_{N}\log(\frac{2r_{N}}{\epsilon})}{2\pi}\int^{\infty}_{0}d\omega\log(1-e^{-\beta_{N}\omega})=\frac{\pi r_{N}\log(\frac{2r_{N}}{\epsilon})}{12\beta_{N}}=\frac{\log(\frac{2r_{N}}{\epsilon})}{24}$ (2.39)
Correspondingly the thermal entropy becomes
$S^{\Phi}_{out}=\frac{1}{24}\log(\frac{2r_{N}}{\epsilon})$ (2.40)
In section (5) we will explicitly compute and will try to match this entropy using the quantization of our proposed UV free bosonic deformed CFT on a ring. We note that, the Hamiltonian of dilatons with stretched horizon boundary condition defined by some UV theory, should be identified with the total Hamiltonian of the observer plus dilatons and the observer clock is set by the ‘time’ generated by static patch Hamiltonian due to Hamiltonian constraint.
3 Dimensional reduction to 2D de-Sitter from Einstein gravity in pure dSD≥3
Unlike higher dimension $(D\geq 4)$, an effective theory in 2D de-Sitter can also be obtained from Einstein gravity with positive cosmological constant in three dimension [40],[29]. For such reduction, one need to consider pure de-Sitter metrics in 3D. To obtain the effective theory, we start with the following ansatz of the 3D metrics [29]:
$ds^{2}=g_{\mu\nu}dx^{\mu}dx^{\nu}+L_{3}^{2}\phi^{2}(x)d\Omega^{2}_{1}$ (3.1)
The Einstein-Hilbert action is
$I_{3}=\frac{1}{16\pi G_{3}}\int_{M}d^{3}X\sqrt{-\hat{g}_{3}}[\hat{R}_{4}-2\Lambda_{3}],\;\Lambda_{3}=\frac{1}{L_{3}^{2}}$ (3.2)
Since $\hat{R}_{3}=R_{2}-\frac{2}{\phi}\nabla^{2}\phi$, we end up with the usual JT gravity up to a total derivative term
$I_{3}=\frac{1}{16\pi G_{2}}\int_{g}d^{2}x\sqrt{-g}[\phi R-2\Lambda\phi-2\nabla^{2}\phi],\;\frac{1}{G_{2}}\equiv\frac{2\pi L_{3}}{G_{3}}$ (3.3)
If one start with the solution of Einstein equation in 3D as pure dS3 in static patch coordinate:
$ds^{2}=-\left(1-\frac{r^{2}}{L_{3}^{2}}\right)dt^{2}+\frac{dr^{2}}{\left(1-\frac{r^{2}}{L^{2}_{3}}\right)}+r^{2}d\theta^{2},$ (3.4)
one should end up with the JT gravity(upto total derivative term) in dS2 static patch:
$ds^{2}=-\left(1-\frac{r^{2}}{L_{3}^{2}}\right)dt^{2}+\frac{dr^{2}}{\left(1-\frac{r^{2}}{L^{2}_{3}}\right)}$ (3.5)
Here $\phi=\frac{r}{L_{3}}$. It is noted that [29], for such pure de-Sitter reduction, $\phi$ must be positive which ensure the positivity of effective Newton’s constant and hence $0<r<L_{3}$. Thus, only half of the full de-Sitter, denoted as the static patch accessible to the observers located on two poles, is described by this metric. This geometry is described by a different Penrose diagram from that of the static patch reduced from higher dimensional near Nariai solution, as noted in [29].
One now observe that in the absence of a GHY term, the sole contribution to the gravitational path integral for JT-dilaton gravity should come from the total derivative or ‘boundary term’. In terms of the angular coordinate $r=L_{3}\cos\theta$, the static patch covers $\frac{\pi}{2}<\theta<\frac{\pi}{2}$. From the Penrose diagram these two are the south and north pole[29]. Hence choosing a suitable ‘boundary’ condition on these poles should determine the static patch entropy due to the presence of cosmological horizon. However, poles are not special points and hence can not be treated as ‘boundary’ of the space-time. Thus choosing some fixed values of dilatons(or it’s derivatives) at those points are strictly ill-defined. At most, one can impose finiteness of the solutions at those points coming from self-adjointness of strum-Liouville type differential operator [^11] See [41] (sec 8.2.1) for a related discussion in the context of solving Laplace equation in spherical polar coordinates.. We will now perform an analogous Weyl rescaling of 2D metrics as we did in the previous section, to rewrite the effective action in a useful form motivated from the previous section.
Using a Weyl scaling $g_{\mu\nu}\rightarrow\omega^{2}(x)\bar{g}_{\mu\nu}$, (3.3) reduces to the following:
$I_{3}=\frac{1}{16\pi G_{2}}\int d^{2}x\sqrt{-\bar{g}_{2}}\left[\phi\bar{R}_{2}+\frac{2}{\omega}(\nabla^{\mu}\phi)(\nabla_{\mu}\omega)-2\Lambda\phi\omega^{2}-2\nabla^{2}\phi-2\nabla^{\mu}\left(\frac{\phi}{\omega}\nabla_{\mu}\omega\right)\right]$ (3.6)
However choosing a similar $\omega$ as in the previous section, we can not recover free field action. In stead choosing $\omega=ce^{\frac{\phi}{4}}$, we end up with the following form:
$I_{3}=\frac{1}{16\pi G_{2}}\int d^{2}x\sqrt{-\bar{g}_{2}}\left[\frac{(\nabla\phi)^{2}}{2}+\bar{R}_{2}\phi-2c\Lambda\phi e^{\frac{\phi}{2}}-2\nabla^{2}\phi-\frac{1}{2}\nabla^{\mu}\left(\phi\nabla_{\mu}\phi\right)\right]$ (3.7)
This type of Liouville-type action (upto total derivative terms) was first observed by Solodukhin in [42] in the context of emergent near horizon conformal symmetry. The potential of the form $\phi e^{\frac{\phi}{2}}$ makes the theory non-conformal which can be thought of as a deformed Liouville action. By studying the equation of motion one can check non vanishing trace of stress tensor [^12] One should note that, the Liouville CFT on a curved manifold can acquire non vanishing constant trace of stress tensor $T=Tr(T_{\mu\nu})$, which is proportional to $\bar{R}_{2}$. However, for this deformed Liouville, apart from the trace anomaly term, $T$ also accompanies with terms involving $\phi$, which makes it strictly non-conformal. as in [42]. Again we can choose $\bar{g}_{\mu\nu}$ to be locally dS2 without loss of any generality. We consider $\bar{g}$ to be the Milne patch, related to static patch by analytic continuation [40], having the following form of metric [43]:
$ds^{2}_{Milne}$ $=-dT^{2}+\sinh^{2}(\frac{T}{L_{3}})d\theta^{2},\;(\text{analytic continuation:}\;r=L_{3}\cosh(\frac{T}{L}),t=\theta)$ (3.8)
$=\frac{1}{\sinh^{2}(\frac{s}{L_{3}})}(-ds^{2}+d\theta^{2}),\;\;\text{using}\;\cosh(\frac{T}{L_{3}})=\coth(\frac{s}{L_{3}})$ (3.9)
$=-\frac{d\hat{t}^{2}}{\hat{t}^{2}-L_{3}^{2}}+(\hat{t}^{2}-L_{3}^{2})d\theta^{2},\;\text{using}\;\coth(\frac{s}{L_{3}})=\frac{\hat{t}}{L_{3}}$ (3.10)
Here the range of time in these coordinate systems are
$-\infty<T<0,\;0<s<\infty,\;L_{3}<\hat{t}$$\infty$ (3.11)
The past conformal boundary corresponds to $T\rightarrow-\infty,s\rightarrow 0,\hat{t}\rightarrow\infty$. Hence in this patch, the late time $s\rightarrow\infty$ should correspond to $\hat{t}\rightarrow L_{3}$. This is the past horizon in static patch. In this late time limit for the past Milne patch, the metric becomes
$ds^{2}_{Milne}|_{s\rightarrow\infty}\approx e^{-\frac{2s}{L_{3}}}(-ds^{2}+d\theta^{2})$ (3.12)
Now the effective action in past Milne patch becomes:
$I_{3}=\frac{1}{16\pi G_{2}}\int dsd\theta\left[-\frac{1}{2}(\partial_{s}\phi)^{2}+\frac{1}{2}(\partial_{\theta}\phi)^{2}+\frac{2\phi}{L_{3}^{2}\sinh^{2}(\frac{s}{L})}(1-ce^{\frac{\phi}{2}})+\text{Total Derivative}\right]$ (3.13)
At past conformal boundary $s\rightarrow 0$, the potential term should dominate over the kinetic term. However, in the late time limit $s\rightarrow\infty$ of past Milne wedge with a time band cut-off $\Lambda_{IR}<s<\infty$, the action becomes a free massless scalar field theory(upto total derivative terms) in the approximate geometry (3.12). Hence the classical solution of $\phi$ can be written as $\phi=\int dkT_{k}(s)e^{ik\theta}$, where $T_{k}(s)=a_{1}e^{iks}+a_{2}e^{-iks}$. The exact solution is determined by imposing the boundary condition at $(s=\Lambda_{IR},\infty)$. The boundary condition should be an UV input and from the classical action we do not have any freedom to choose a meaningful one. Since specifying boundary condition on future slice would affect the initial boundary condition, it can make a potential violation of causality[44]. The correct way is to specify initial condition and evolve it with the correct evolutionary operator [^13] That necessarily would be an UV input. capturing the full dynamics; which could further tell us the correct ‘boundary condition’ on the approximate time band.
We end this section by making two comments:
  • The problem of imposing unphysical ‘boundary’ condition on poles in the static patch can be controlled by going into the past Milne patch and by using late time effective conformal theory as we explained above. However, it is still unclear about the correct physical boundary condition one should impose at the boundary of the time band purely from this classical action. We note that, the similar analysis can be carried out for higher dimensional reduction of pure de-Sitter geometries. Following the similar way [^14] The details are not present in this present note to avoid unnecessary cluttering., one could obtain an effective theory of the past Milne patch with a different potential term. Interestingly, one again obtains an universal feature showing the early time dominance of potential term(of different form for 2d reduction from different dimensional Einstein gravity) and the late time dominance of free bosonic CFT.
  • The effective dynamical theory in the past Milne patch of being non conformal to conformal one, can be regarded as some kind of time dependent RG flow. More intuitively, one may map this into certain quench problem, where initial gapped theory is evolved by a gapless CFT Hamiltonian after certain Lorentzian time ($\Lambda_{IR}$). This will motivate us to study certain deformed CFT Hamiltonian in the next section, which breaks time translation symmetry upon acting on an initial vacuum state of a gapped theory at the past infinity and describes time dependent dynamics of the Milne patch.
4 Deformed CFT, Milne patch and critical quench
Effectively, to capture the effective Milne patch dynamics(full) obtained in the past section, one need to find a time dependent evolution operator that is best described in a smooth critical quench set-up[45]-[49]. Given the Lagrangian of the effective theory, it seems analytically less tractable in such a set-up. However, in a crude sense, we can still approximate the dynamics as an instantaneous critical quench protocol [50], which enables us to study post quench dynamics inside the approximate time band $s:(\Lambda_{IR},\infty)$. In [51], Cardy and Calabrese pioneered an analytically tractable critical quench set-up, where the initial state is explicitly written down as $|\psi\rangle_{in}=e^{-\tau_{0}H_{CFT}}|B\rangle$, which evolves to a late time steady state under post quench CFT evolution [52]. Here $\tau_{0}$ encodes the mass-gap parameter of the initial gapped Hamiltonian. $|B\rangle$ refers to the conformal boundary state. The justification of choosing such state comes from the RG sense where pre-quench and post-quench Hamiltonian is related by small irrelevant deformation under which any boundary condition flows to a conformal boundary condition. A relatively nicer argument was presented in [53] [^15] We thank Bobby Ezhuthachan for explaining this., where the Euclidean time strip picture is more naturally appeared from analytic continuation of Euclidean correlators to Lorentzian one via $i\epsilon$ prescription. To mention it briefly, consider the correlation function
$\langle\psi|e^{iHt_{1}+\epsilon_{1}H}\mathcal{O}(x_{1})e^{-iHt_{12}-\epsilon_{12}H}\mathcal{O}(x_{2})e^{-iHt_{23}-\epsilon_{23}H}\dots\mathcal{O}(x_{n})e^{-iHt_{n}-\epsilon_{n}H}|\psi\rangle$ (4.1)
Here we used the analytic continuation $\tau\rightarrow it+\epsilon$ and the notation $t_{ij}=t_{i}-t_{j}$. According to the $i\epsilon$ prescription, this correlation function is analytic when $(0>\epsilon_{1}>\epsilon_{2}>\dots>\epsilon_{n}>0)$, which does not make any sense. However this would not be a problem when $|\psi\rangle$ is an eigenstate of $H$. The problem occurs only when the initial state is not an eigenstate of the Hamiltonian. The effective way to get rid of this problem, is to formally introduce a cut-off $\tau_{0}$ to regulate high energy modes ($\frac{1}{\tau_{0}}$). Hence the regulated state should be $e^{-\tau_{0}H}|\psi\rangle$, for which the analytic continuation has the following order: $\tau_{0}>\epsilon_{1}>\epsilon_{2}>\dots>\epsilon_{n}>-\tau_{0}$. In this Euclidean picture, the correlation function is computed in a strip of Euclidean time width $2\tau_{0}$. The boundary condition at the edges of this strip is chosen to be conformal one following the CC ansatz.
To map this set-up with the effective Milne patch quench, we may think $\tau_{0}$ to be certain scale dictated by the non conformal initial theory. Here cleverly, the problem of imposing boundary condition in the effective Lorentzian time width is replaced by imposing conformal boundary condition at the edges of the strip in Euclidean time. This no longer violates any causal requirement of the original Lorentzian problem. Nevertheless, the Hamiltonian of the CFT is still needed to be the same of that of a free massless bosons in the approximate geometry (3.12) of late time band in past Milne wedge. This will motivate us to find a deformed CFT Hamiltonian, that captures the dynamics of Milne wedge as well as mimics the set-up of the critical sudden quench.
4.1 Boundary non-preserving deformed CFT Hamiltonian on a strip and Milne patch dynamics
[…]
$ds^{2}=4L^{2}_{AdS}\frac{dzd\bar{z}}{(z-\bar{z})^{2}}=4L^{2}_{AdS}\left(\frac{\partial z}{\partial\omega}\right)\left(\frac{\partial\bar{z}}{\partial\bar{\omega}}\right)\frac{d\omega d\bar{\omega}}{(z(\omega)-\bar{z}(\bar{\omega}))^{2}}\rightarrow 4L^{2}_{AdS}\mu^{2}\frac{ds_{E}^{2}+d\theta^{2}}{\sin^{2}2\mu s_{E}}$ (4.2)
Here $z¯=tix\bar{z}=t-ix$ and $\omega=\theta+is_{E}$, $\bar{\omega}=\theta-is_{E}$. After analytically continuing $s_{E}\rightarrow is$ and $L_{AdS}\rightarrow iL$ [^16] Here $L$ is the de-Sitter radius. with setting $2\mu L=1$, one could end up with the standard Milne wedge metric described in the previous section:
$ds^{2}=4L^{2}\mu^{2}\frac{-ds^{2}+d\theta^{2}}{\sinh^{2}{2\mu s}}=\frac{-ds^{2}+d\theta^{2}}{\sinh^{2}{\frac{s}{L}}}$ (4.3)
[…]
$H_{o}=\frac{\gamma}{\pi}\int^{\pi}_{0}du\;\sin(u)T_{00}(u),\;\gamma\in\mathbb{R}$ (4.4)
After mapping to Upper Half Plane (UHP) using the map ($w\rightarrow z=e^{w}=t+ix$), we end up with the Hamiltonian on the UHP:
$H=i\gamma(L_{1}-L_{-1})-i\gamma(\bar{L}_{1}-\bar{L}_{-1})$ (4.5)
Here $L_{n},\bar{L}_{n}$ are defined as
$L_{n}=\frac{c}{24}\delta_{n,0}+\frac{1}{2}\int^{\pi}_{0}\frac{du}{\pi}e^{inu}T(u)=\frac{c}{24}\delta_{n,0}+\frac{1}{2\pi i}\int_{C}dzz^{n+1}T(z)$ (4.6)
$\bar{L}_{n}=\frac{c}{24}\delta_{n,0}+\frac{1}{2}\int^{\pi}_{0}\frac{du}{\pi}e^{-inu}\bar{T}(u)=\frac{c}{24}\delta_{n,0}+\frac{1}{2\pi i}\int_{\bar{C}}d\bar{z}\bar{z}^{n+1}\bar{T}(\bar{z}).$ (4.7)
Here, the contours $C$ and $\bar{C}$ refer to a half circle around the origin of UHP. For details, we refer [1].
The Hamiltonian (4.5) is Hermitian since $L_{n}^{\dagger}=L_{-n}$. However, the same Hamiltonian does not preserve the symmetry of the UHP and the conformal boundary condition. One can simply check it by taking the Euclidean space-time generators $(l_{n}=z^{n+1}\partial_{z},\bar{l}_{n}=\bar{z}^{n+1}\partial_{\bar{z}}$ corresponding to the Virasoro modes $(L_{n},\bar{L}_{n})$ and act on the real boundary of the UHP as:
$i\gamma(l_{1}-l_{-1}-\bar{l}_{1}+\bar{l}_{-1})(z-\bar{z})|_{z=\bar{z}}\neq 0$ (4.8)
[…]
As in [54], to find the proper conformal frame of the Euclidean time($s_{E}$) evolution generated by $H_{o}$, we need to find the curves $(z(s),\bar{z}(s))$ solving the curve equation:
$i\gamma(l_{1}-l_{-1}-\bar{l}_{1}+\bar{l}_{-1})=\frac{\partial z}{\partial s}\frac{\partial}{\partial z}+\frac{\partial\bar{z}}{\partial s}\frac{\partial}{\partial\bar{z}}$ (4.9)
This involves solving
$\int\frac{dz}{i\gamma(z^{2}-1)}=\int ds_{E}$ (4.10)
Similar integral equation exists for $\bar{z}$ with $i\rightarrow-i$. Solving this, we get the following curves:
$z=-\coth\gamma(\theta+is_{E})\equiv-\coth{\gamma\omega},\;\bar{z}=-\coth{\gamma(\theta-is_{E})}\equiv-\coth{\gamma\bar{\omega}}.$ (4.11)
Here $\theta$ is an integration constant of the curve equation. The boundary of the UHP $z=\bar{z}$ maps to the following:
$z=\bar{z}\implies s_{E}=(0,\frac{\pi}{\gamma})$ (4.12)
Hence the effective geometric description of $(\omega,\bar{\omega})$ coordinates is an Euclidean time-strip of width $\frac{\pi}{\gamma}$ and $\theta:(-\infty,\infty)$. To find the constant $s_{E}$ and constant $\theta$ curve in the UHP, we consider the following:
$\frac{z-1}{z+1}=e^{2\gamma(\theta+is_{E})},\;\frac{\bar{z}-1}{\bar{z}+1}=e^{2\gamma(\theta-is_{E})}$ (4.13)
From this, to study the constant $\theta$ curve or the evolution of $s_{E}$, we would consider the following:
$e^{4\gamma\theta}=\frac{(z-1)(\bar{z}-1)}{(z+1)(\bar{z}+1)}$ (4.14)
This can be simplified to the following equation:
$(t+\coth{2\gamma\theta})^{2}+x^{2}=(\operatorname{csch}2\gamma\theta)^{2}$ (4.15)
Since $x\geq 0$, the constant $\theta$ curves generates semicircles of radius $\operatorname{csch}2\gamma\theta$ centered around $(t=-\coth{2\gamma\theta},\;x=0)$. Thus the range of $s_{E}:(0,\pi/\gamma)$. At $\theta\rightarrow\pm\infty$, the curves shrinks to the point $t=\pm 1$. Consequently, to find the $\theta$ curves of constant $s_{E}$ curves, we consider the following:
$e^{4i\gamma s_{E}}=\frac{(z-1)(\bar{z}+1)}{(z+1)(\bar{z}-1)}$ (4.16)
After some algebraic manipulation, we end up with the following curve equation:
$t^{2}+(x-\cot{2\gamma s_{E}})^{2}=(\csc 2\gamma s_{E})^{2}$ (4.17)
[…]
Figure 1: Curves of $\gamma\theta$ and $\gamma s_{E}$ on the UHP($t,x$)
One can see that $t=\pm 1$ and $x=0$ is the trivial solution of the curve for all $s_{E}$. Hence $(t=\pm 1,\;x=0)$ are the fixed point of the Hamiltonian flow. Hence constant $s_{E}$ or the $\theta$ curves are semicircles end on the fixed points. This translates the fixed points to be $\theta=\pm\infty$. The curves are shown in the fig(1).
One can also obtain the fixed points by solving $z(s_{E})=z(0)$. Here $z(s)=\coth{\gamma}(\theta+is_{E})$ and $z(0)=\coth{\gamma\theta}$. From this we get
$z(s)=\frac{z(0)\cos{\gamma s_{E}}+i\sin{\gamma s_{E}}}{iz(0)\sin{\gamma s_{E}}+\cos{\gamma s_{E}}}\equiv\frac{az(0)+b}{cz(0)+d}\;\text{with}\;ad-bc=1$ (4.18)
Hence, the flow is mapped to a Mobius transformation with imaginary $b,c$. The solution of $z(s)=z(0)$ is $z=\pm 1$. This is the fixed point we got earlier which corresponds to $\theta=\pm\infty$. We note that this type of Hamiltonian flow has been studied in the context of Floquet CFTs [55],[56]. In particular, [56] studied a Floquet drive for which the dynamical heating phase has similar structure of Mobius transformation [^18] In the same paper [56], non-trivial dynamics in entanglement asymmetry has been found for a similar driven CFT on a strip. The initial global symmetry is broken by the boundary and the boundary non-preserving Floquet drive would never restore the symmetry. This result and related observation provides one of the major motivations behind this present note..
Following the observation of [1], we should consider the metric of UHP as the AdS2 Poincare. Hence the effective metric under such flow becomes the Euclidean Milne wedge in de-Sitter as we described in (4.2). Upon analytic continuation to $s_{E}\rightarrow is$, we will get Lorentzian Milne patch with $s:(0,\infty)$. To study the time dependence, one may consider one-point function of CFT scalar primary of dimension $h$ in $(\omega,\bar{\omega})$ plane, by mapping it to UHP.
$\langle\mathcal{O}(\omega,\bar{\omega})\rangle=\left(\frac{\partial z}{\partial\omega}\right)^{h}\left(\frac{\partial\bar{z}}{\partial\bar{\omega}}\right)^{h}\langle\mathcal{O}(z,\bar{z})\rangle_{UHP}=\left(\frac{a_{\mathcal{O}}\gamma}{i\sin{2\gamma s_{E}}}\right)^{2h}$ (4.19)
Here $a_{\mathcal{O}}$ is an undetermined constant. By doing $s_{E}\rightarrow is$, we get
$\langle\mathcal{O}(s,\theta)\rangle=\left(\frac{\gamma}{\sinh{2\gamma s}}\right)^{2h}$ (4.20)
Note that the late (real)time behavior coincides with the one point function under quantum critical quench [52], which decays exponentially at late time $s>>\frac{1}{2\gamma}$.
However, unlike the CC ansatz, here we took slightly different initial state. Our initial state has the following form:
$|\psi\rangle_{in}=e^{-H_{o}}|0\rangle_{S}$ (4.21)
[…]
4.2 Emergence of steady state at late time
We have now seen the deformed CFT on the strip has profound similarity with global quench set-up followed by CC ansatz. To be more precise, the current set-up should correspond to an inhomogeneous critical quench[57],[53]. Here $1/\gamma$ plays the role of the cut-off $\tau_{0}$. However a subtle difference arises in the energy momentum tensor. While, in the CC state, the $\langle T\rangle_{CC}=\frac{\pi c}{24(2\tau_{0})^{2}}>0$, in our prescribed non-conformal Milne vacuum:
${}_{in}\langle\psi|T(\omega)|\psi\rangle_{in}=-\frac{c\gamma^{2}}{12}<0$ (4.22)
Since we are interested in the physical $\gamma\rightarrow\infty$ limit, the stress energy becomes negatively divergent. This would further indicates the state $|\psi\rangle_{in}$ is unstable. To make further support on this, we will now compute two point primary correlation function on a fixed time slice $s$, following [52],[58]:
${}_{in}\langle\psi|\mathcal{O}(s_{E},0)\mathcal{O}(s_{E},\theta)|\psi\rangle_{in}$
$=\left(\frac{\partial z_{1}}{\partial\omega_{1}}\right)^{h}\left(\frac{\partial\bar{z}_{1}}{\partial\bar{\omega}_{1}}\right)^{h}\left(\frac{\partial z_{2}}{\partial\omega_{2}}\right)^{h}\left(\frac{\partial\bar{z}_{2}}{\partial\bar{\omega}_{2}}\right)^{h}\langle\mathcal{O}(z_{1},\bar{z}_{1})\mathcal{O}(z_{2},\bar{z}_{2})\rangle_{UHP}$ (4.23)
Here $z_{1,2}=\coth{\gamma\omega_{1,2}}$, $\bar{z}_{1,2}=\coth{\gamma\bar{\omega}_{1,2}}$ and
$\omega_{1}=is_{E},\bar{\omega}_{1}=-is_{E};\;\omega_{2}=\theta+is_{E},\bar{\omega}_{2}=\theta-is_{E}.$ (4.24)
The UHP two point correlator has the following structure
$\langle\mathcal{O}(z_{1},\bar{z}_{1})\mathcal{O}(z_{2},\bar{z}_{2})\rangle_{UHP}=\left(\frac{z_{12}z_{2\bar{1}}}{z_{12}z_{\bar{1}\bar{2}}z_{1\bar{1}}z_{2\bar{2}}}\right)^{2h}F(\eta)$ (4.25)
Here the notation $z_{ij}=z_{i}-z_{j}$ and $F(\eta)$ is an undetermined function of cross ratio $\eta$. For free scalar CFT, one can exactly compute this two point function. However, here we consider general CFT to find universal late time behavior. Using (4.24), we find
$\eta=\frac{\sin^{2}{(2\gamma s_{E})}}{\sinh{\gamma(\theta-2is_{E})}\sinh{\gamma(\theta+2is_{E})}}$ (4.26)
After analytic continuation $s_{E}\rightarrow is$, we get
$\eta=\frac{1-\cosh{4\gamma s}}{\cosh{2\gamma\theta}-\cosh{4\gamma s}}$ (4.27)
We now consider the late time limit as well as $s,\theta>>\frac{1}{4\gamma}$ [^20] We are interested only in this limit, since the effective dilaton theory becomes approximately conformal only at late time.. We can notice two following different behavior of $\eta$:
$\eta\sim 1,\;\text{when}\;s>\frac{\theta}{2}$ (4.28)
$\eta\sim e^{4\gamma(s-\frac{\theta}{2})}\sim 0\;\text{when}\;s<\frac{\theta}{2}$ (4.29)
We now compute
$\left(\frac{z_{12}z_{2\bar{1}}}{z_{12}z_{\bar{1}\bar{2}}z_{1\bar{1}}z_{2\bar{2}}}\right)^{2h}=\gamma^{4h}\left[\frac{\cosh{2\gamma\theta}-\cosh{4\gamma s}}{\sinh^{2}{\gamma\theta}\sinh^{2}{2\gamma s}}\right]^{2h}$ (4.30)
Again considering $s,\theta>>\frac{1}{4\gamma}$, we find
$\left(\frac{z_{12}z_{2\bar{1}}}{z_{12}z_{\bar{1}\bar{2}}z_{1\bar{1}}z_{2\bar{2}}}\right)^{2h}\approx\Bigg\{\begin{split}(\gamma)^{4h}e^{-4\gamma h\theta}\;\text{for}\;s>\frac{\theta}{2}\\ (\gamma)^{4h}e^{-8\gamma hs}\;\text{for}\;s<\frac{\theta}{2}\end{split}$ (4.31)
While $F(1)=1$, at $\eta\rightarrow 0$, $F(\eta)\sim(a_{\mathcal{O}})^{2}\eta^{2h_{b}}$. Here, $h_{b}$ is the boundary operator dimension which appears in the leading term of the bulk-boundary OPE of $\mathcal{O}(z,\bar{z})$. Finally putting all these results, we obtain the following expression for two point function at late time limit:
${}_{in}\langle\psi|\mathcal{O}(s_{E},0)\mathcal{O}(s_{E},\theta)|\psi\rangle_{in}\sim\Bigg\{\begin{split}&\gamma^{4h}e^{-4\gamma h\theta}\;\text{for}\;s>\frac{\theta}{2}\\ &\gamma^{4h}a_{\mathcal{O}}^{2}e^{8\gamma h_{b}(s-\frac{\theta}{2})}e^{-8\gamma hs}\;\text{for}\;s<\frac{\theta}{2}\end{split}$ (4.32)
Hence we can see the late time behavior of the two point correlator for $s>\theta/2$ becomes constant, while for $s<\theta/2$ it is decaying exponentially. Similarly, one can compute entanglement entropy by twist operator two point function on a constant $s$ slice and get  [^21] In the following expression #\# refers to some constant.
$S_{EE}(s)\sim\Bigg\{\begin{split}&\frac{c}{3}\ln\frac{1}{2\gamma}+(\#).s\;\text{for}\;s<\frac{\theta}{2}\\ &\frac{c}{3}\ln\frac{1}{2\gamma}+(\#).\theta\;\text{for}\;s>\frac{\theta}{2}\end{split}$ (4.33)
Hence again the late time entanglement entropy shows steady state values as in the global quench scenario[52]. In this way, we can see the effective dilaton dynamics in the (past) Milne wedge, governed by the proposed Milne vacuum $|\psi\rangle_{in}$ is fundamentally unstable and approaches to a steady state at a late but finite time. This resembles with the universal feature of global quantum critical quench.
5 Deformed CFT, static patch and observer’s entropy
In Lorentzian time, the static patch and Milne wedge provide the same description [30] with different realization of isometries and coordinates representation related by analytic continuation as in (3.8). We can try to do similar continuation in Euclidean time. For instance, we analytically continue $s_{E}\rightarrow i\theta^{\prime}+\frac{\pi}{2L}$ and $\theta\rightarrow it^{\prime}$ such that
$ds^{2}=-\frac{ds_{E}^{2}+d\theta^{2}}{\sin^{2}{\frac{s_{E}}{L}}}\rightarrow\frac{d\theta^{\prime 2}+dt^{\prime 2}}{\cosh^{2}{\frac{\theta^{\prime}}{L}}}$ (5.1)
Here $\theta^{\prime}:(-\infty,\infty)$. Upon Lorentzian continuation $t^{\prime}\rightarrow i\tau$, we will end up with the familiar static patch metric in conformal coordinates as in (2.29) [^22] We need to identify $L=r_{N}$.:
$ds^{2}=\frac{-d\tau^{2}+d\theta^{2}}{\cosh^{2}{\frac{\theta}{L}}}$ (5.2)
Here we change the notation $\theta^{\prime}$ to $\theta$ for this section to avoid excessive notational clutter.
However doing such analytic continuation $s_{E}\rightarrow s^{\prime}_{E}+\frac{\pi}{2L}$ followed by $s^{\prime}_{E}\rightarrow i\theta$, is not possible in the strip picture. The Euclidean time boundary at $s_{E}=0,\frac{\pi}{L}$ breaks the time translational invariance which prevents such analytic continuation. However, we know in static patch has explicit time translational isometry and the Euclidean time should be periodic to remove the conical singularity at the horizons. This problem can be circumvented if we go from strip to ring or cylinder with the same Hamiltonian $H_{c}$ of the form (4.5) [^23] One may alternatively think of this like joining another strip to make a cylinder with changing the boundary condition from Dirichlet to periodic. Heuristically, this is just adding the upper Milne wedge and then analytically continue to extended static patch.. Since the analytic continuation does not change the map from full complex plane to $(\theta,t^{\prime})$ $z\rightarrow\omega$, the form of the Hamiltonian will not change on the ring. Instead of semicircles, the Euclidean time curve will now become a full circle manifesting time translation symmetry(see fig(2). This motivates us to study deformed CFT with Hamiltonian of the form (4.5) on a ring. For this section, we will use the same notation for $L_{n},\bar{L}_{n}$ which describes Virasoro modes on the full complex plane.
5.1 Deformed CFT on a ring, emergent Virasoro algebra and stretched horizon in static patch
The deformed Hamiltonian (4.5) on full complex plane can be written as [^24] Here we ignore the Casimir term appearing due to the conformal mapping from cylinder to plane.
$H_{c}=\frac{i\gamma}{2\pi i}\oint dz\;(z+1)(z-1)T(z)-\frac{i\gamma}{2\pi i}\oint d\bar{z}\;(\bar{z}+1)(\bar{z}-1)\bar{T}(\bar{z})$ (5.3)
Using $z=\coth{\gamma\omega}$ and $\bar{z}=-\coth{\gamma\bar{\omega}}$, with $\omega=\theta+it^{\prime}$$\bar{\omega}=\theta-it^{\prime}$ and the transformation of stress tensor $T(\omega)=(\frac{\partial z}{\partial\omega})^{2}T(z)+\frac{c}{12}Sch\{z,\omega\}$, we end up with:
$H_{c}$ $=\frac{1}{2\pi}\left[\int^{\infty}_{-\infty}d\theta\;T(\omega)-\int^{-\infty}_{\infty}d\theta\;\bar{T}(\bar{\omega})-\frac{c}{12}\left(\int^{\infty}_{-\infty}d\theta\;Sch\{z,\omega\}-\int^{-\infty}_{\infty}d\theta\;Sch\{\bar{z},\bar{\omega}\}\right)\right]$
$=\frac{1}{2\pi}\int^{\infty}_{-\infty}d\theta\;(T(\omega)+\bar{T}(\bar{\omega}))+\frac{c\gamma^{2}}{6\pi}\int^{\infty}_{-\infty}d\theta$ (5.4)
Here we have used constant $t^{\prime}$ curve as the contour of integration upon mapping from constant time circles in the complex plane. We can see the Schwarzian term will be diverging and we need to cut-off the $\theta$ integral as in [1]. We choose $\gamma\theta:(-\Lambda_{c},\Lambda_{c})$. The need for cut-off also arises to define Virasoro modes as in modular quantization [59], [28],[60]. After imposing the cut-off the constant term looks similar to a constant Schwarzian term appeared as an ADM mass term of JT gravity with negative cosmological constant. As we mentioned in the introduction 1, the ADM clock located at asymptotic boundary plays the role of observer. However, the appearance of this constant Schwarzian term has no clear interpretation to us without a GHY boundary term.
We should also mention that the vacuum of the static patch can be identified with the conformal vacuum on cylinder $|0\rangle_{c}$ such that $H_{c}|0\rangle_{c}=0$. In this vacuum, by using conformal map $z\rightarrow\omega$ and analytic continuation from $t^{\prime}\rightarrow\tau$, one can obtain two point function $_{c}\langle 0|\mathcal{O}(\tau,0)\mathcal{O}(0,0)|0\rangle_{c}\propto(\sinh\gamma\tau)^{-2\Delta}$, which shows thermal nature.
[…]
$\partial_{t^{\prime}}H_{c}=\frac{1}{2\pi}\left[T(\infty)-T(-\infty)-\bar{T}(\infty)+\bar{T}(-\infty)\right]$ (5.5)
Now $T(\omega)=\left[\frac{1}{\sinh^{4}{\gamma\omega}}T(z)+\frac{c\gamma^{2}}{12\pi}\right]$ and $\bar{T}(\bar{\omega})=\left[\frac{1}{\sinh^{4}{\gamma\bar{\omega}}}T(z)+\frac{c\gamma^{2}}{12\pi}\right]$. Since $\theta\rightarrow\pm\infty$ corresponds to $z\rightarrow\pm 1$, $T(z)|_{z=\pm 1}$, $\bar{T}(\bar{z})|_{\bar{z}=\pm 1}$ is finite. Thus we can immediately see [^25] Here again the stress tensor is negative but finite. Even though, one can have thermal two point function and other thermal features of the static patch, this constant negative stress tensor is not concerning. Rather the difference from an positive stress tensor in the AdS2 black hole vacuum is quite encouraging in the sense of distinguished nature of a cosmological horizon and a black hole horizon. An observer inside the cosmological horizon can effectively measure negative local energy density. However this static patch vacuum is not a (full) de-Sitter invariant vacuum. Time translation invariance in an expanding universe is just an artifact of forming causal horizons enclosing the observers.
$T(\omega)|_{\theta=\pm\infty}=\bar{T}(\bar{\omega})|_{\theta=\pm\infty}=-\frac{c\gamma^{2}}{12}$ (5.6)
Hence, we have $T(\omega)-\bar{T}(\bar{\omega})]|_{\theta=\pm\infty}=0$, that manifests $\partial_{t^{\prime}}H_{c}=0$. This is what we dubbed as emergent conformal boundary condition at the horizons $\theta=\pm\infty$. We should emphasize that the fixed points of the deformed Hamiltonian are realized as emergent conformal boundary in the proper time frame of static patch.
For a free bosonic CFT, the emergent horizon boundary condition can be either a Dirichlet one or Neumann one. Both will give the similar solutionson field modes as we discussed in section 2. After imposing the stretched horizon cut-off we can impose the same condition on the regulator.
We can now construct emergent Virasoro modes as first discussed in the context of dipolar quantization [61],[62] [^26] See also [63], for a Lorentzian analogue.. Following [62], we define
$H_{c}=\mathcal{L}_{0}+\bar{\mathcal{L}}_{0}$ (5.7)
The corresponding infinitesimal generators are
$\tilde{l}_{0}=i\gamma(z^{2}-1)\partial_{z},\;\bar{\tilde{l}}_{0}=-i\gamma(\bar{z}^{2}-1)\partial_{\bar{z}}$ (5.8)
Hence the eigenfunction of $\tilde{l}_{0}$ is $f_{k}$ with eigenvalue $k$.
$f_{k}(z)=e^{k\int_{z}\frac{dz}{i\gamma(z^{2}-1)}}=\left(\frac{z-1}{z+1}\right)^{-\frac{ik}{2\gamma}}$ (5.9)
Thus we can find the infinitesimal modes $\tilde{l}_{k}=i\gamma(z^{2}-1)f_{k}(z)$ and the corresponding conserved charges defined as
$\mathcal{L}_{k}=\frac{\gamma}{2\pi}\int_{C}dz\;(z-1)^{-\frac{ik}{2\gamma}+1}(z+1)^{\frac{ik}{2\gamma}+1}T(z)$ (5.10)
Here $C$ refers to the constant $t^{\prime}$ contour which denotes a $\theta$ integral over $(\infty,-\infty)$. To construct the commutators of $\mathcal{L}_{k}$, one again need to specify a cut-off at $z=\pm 1+\epsilon e^{i\theta_{\epsilon}}$.
[…]
Figure 2: Curves of $\theta$ and $t^{\prime}$ on the full complex plane
Since the contour of the $(t^{\prime}-\theta)$ as in fig(2), is just the $\pi/2$ rotated version of the contour of modular quantization in cylinder [28], one can check the end result would be similar to that one. We avoid the repetitive computation and we refer to [28],[1] for the details. We finally obtain:
$[\mathcal{L}_{k}^{\Lambda_{c}},\mathcal{L}_{k^{\prime}}^{\Lambda_{c}}]=(k-k^{\prime})\mathcal{L}_{k+k^{\prime}}^{\Lambda_{c}}+\frac{c_{eff}}{12}(k^{3}+4\gamma^{2}k)\delta_{k+k^{\prime}}$ (5.11)
$\mathcal{L}_{k}^{\Lambda_{c}}$ is the modified generator having the contour length of $2\Lambda_{c}\equiv\frac{2}{\gamma}\log\left(\frac{2}{\epsilon}\right)$ and $\theta:(-\Lambda_{c},\Lambda_{c})$. Here $c_{eff}=\frac{c}{2\pi\gamma}\ln\left(\frac{2}{\epsilon}\right)$. Here again $k$ is discretized and of the similar form for stretched horizon normal modes obtained in (2.32):
$k=\frac{2n\pi\gamma}{\ln\left(\frac{2}{\epsilon}\right)}$ (5.12)
[…]
5.2 Modular quantization on static patch and thermal partition function
To quantize the Hamiltonian we can again use the usual modular quantization technique developed in [28],[1]. Again we are not repeating the procedure. We summarize the key points and some important clarifications from [1]:
  • Due to cut-off with conformal boundary condition, the quantization contour of $(t^{\prime},\theta)$ curves can be mapped to a thermal BCFT on an annulus of width $W=2\pi\Lambda_{c}$.
  • We can construct eigenstates of regulated Hamiltonian $(\mathcal{L}_{0}^{\Lambda_{c}}+\bar{\mathcal{L}}_{0}^{\Lambda_{c}})$ which are localized at the cut-off surface or stretched horizon. These eigenstates are constructed out of primaries as well as modular Virasoro descendants. We note that, these states are emergent due to the emergent modular Virasoro algebra and the conformal boundary condition at the cut-off. All the states at the cut-off surface can be constructed explicitly for massless free bosons as states of regulated Fock space, briefly described in section 2.
  • Using this quantization, the $\epsilon\rightarrow 0$ limit can be identified with the zero modes of the Hamiltonian as we will show now. However, mapping back to $(\omega,\bar{\omega})$ plane these primaries are vanishing, leaving only vacuum at the horizon [^27] We took a different claim for these than our earlier paper [1], which we think is wrong.. Hence the vanishing edge modes should further give a vanishing micorcanonical entropy count sourced purely from the surviving vacuum state [^28] One can not obtain this just naively putting $\epsilon=0$ for which it is diverging. This different results further suggests that measuring entropy without a cut-off with boundary condition is meaningless.. This on the other hand tells, the entropy explicitly measures the stretched horizon observer entropy, which is not obtained in the semiclassical limit without an observer. The same applies in AdS, where one must impose a cutoff on the dilaton, which can be made arbitrarily large but cannot be taken literally to infinity.
Following [28], we consider states $|\tilde{0}\rangle$ of the form
$|\tilde{0}\rangle=\sum_{n}\sum_{m}a_{n}b_{m}L_{-1}^{n}\bar{L}_{-1}^{m}\mathcal{O}_{h}(0,0)|0\rangle_{c}$ (5.13)
We want to construct those satisfying
$H_{c}|\tilde{0}\rangle=0$ (5.14)
Using the identities:
$[L_{1},L^{n}_{-1}]=n(n-1)L^{n-1}_{-1}+2nL^{n-1}_{-1}L_{0}$
$[L_{0},L_{-1}^{n}]=nL^{n}_{-1},$ (5.15)
we obtain
$i\gamma(L_{1}-L_{-1})\sum_{n}a_{n}L_{-1}^{n}\mathcal{O}_{h}(0,0)|0\rangle_{c}$
$=\sum_{n}i\gamma\left(a_{n+1}n(n+1)+2(n+1)a_{n+1}h-a_{n-1}\right)L^{n}_{-1}\mathcal{O}_{h}(0,0)|0\rangle_{c}$ (5.16)
Similarly
$i\gamma(\bar{L}_{1}-\bar{L}_{-1})\sum_{m}b_{n}\bar{L}_{-1}^{m}\mathcal{O}_{h}(0,0)|0\rangle_{c}$
$=\sum_{m}i\gamma\left(a_{m+1}m(m+1)+2(m+1)b_{m+1}h-b_{m-1}\right)\bar{L}^{m}_{-1}\mathcal{O}_{h}(0,0)|0\rangle_{c}$ (5.17)
Hence combining this and choosing $a_{n}=\frac{a^{n}}{n!}$, $b_{m}=\frac{b^{m}}{m!}$, we have
$\left[i\gamma(L_{1}-L_{-1})-i\gamma(\bar{L}_{1}-\bar{L}_{-1})\right]|\tilde{0}\rangle$
$=i\gamma\sum_{n,m}\frac{a^{n}b^{m}}{n!m!}\left[an+2ah-\frac{n}{a}-bm-2bh+\frac{m}{b}\right]L^{n}_{-1}\bar{L}^{m}_{-1}\mathcal{O}_{h}(0,0)|0\rangle_{c}$ (5.18)
We can readily see that when $a=b=1$ or a=b=1a=b=-1, we get the desired result i.e. $H_{c}|\tilde{0}\rangle=0$. Hence the zero-mode vacuum state becomes
$|\tilde{0}\rangle=\Bigg\{\begin{split}&e^{L_{-1}}e^{\bar{L}_{-1}}\mathcal{O}_{h}(0,0)|0\rangle_{c}=\mathcal{O}_{h}(1,1)|0\rangle_{c}\\ &e^{-L_{-1}}e^{-\bar{L}_{-1}}\mathcal{O}_{h}(0,0)|0\rangle_{c}=\mathcal{O}_{h}(-1,-1)|0\rangle_{c}\end{split}$ (5.19)
These are precisely the primaries $\mathcal{O}(z,\bar{z})$ located at the fixed points $z,\bar{z}=\pm 1$. However, mapping into the $\omega,\bar{\omega}$ plane, we can see $\mathcal{O}(\omega,\bar{\omega})|_{\theta=\pm\infty}$ is vanishing due the Jacobian factor of the conformal mapping for primaries.
$\left[\left(\frac{dz}{d\omega}\right)^{h}\left(\frac{d\bar{z}}{d\bar{\omega}}\right)^{h}\right]_{\theta=\pm\infty}=0$ (5.20)
Hence only vacuum or identity operator survives exactly at the $\epsilon\rightarrow 0$ limit. This again shows the Hilbert space of modular quantization is completely different and emergent due to the boundary condition at near horizon cut-off.
To compute the thermal entropy of the annulus we again follow the closed-open string duality [64]. The annulus partition function of temperature $\frac{4\pi^{2}}{W}$ can be obtained using closed string channel partition function in the annulus of length $W=2\nu\Lambda_{c}=2\log(\frac{2}{\epsilon})$ as
$Z_{\epsilon}={}_{\epsilon}\langle B|e^{-W(L_{0}+\bar{L}_{0}-\frac{c}{12})}|B\rangle_{\epsilon}$ (5.21)
In $\epsilon\rightarrow 0$, similarly to [28], we end up with
$S_{\epsilon\rightarrow 0}=\frac{c}{3}\log(\frac{2}{\epsilon})+g_{a}+g_{b}$ (5.22)
Hence the leading term matches with the entanglement entropy in a full time slice of the static patch $\theta:(-\Lambda_{c},\Lambda_{c})$ as one can check by computing twist operator two point function using the map $z\rightarrow\omega$. This is again consistent with the fact that the time translation of static patch Hamiltonian is identified with the flow by modular Hamiltonian on the full patch bounded by the horizons.
On the other hand, we can compare this result by explicit computation of free massless bosons($c=1$) in the stretched horizon background as in (2.40). We have seen a potential mismatch of overall factor of $\frac{1}{8}$. If we have 8 massless scalars in the static patch background, this would exactly match with a single scalar in the worldsheet description. Currently we have no clear understanding behind this mismatch. Neglecting that potential mismatch, we can see both of these terms proportional to $\log(\frac{2}{\epsilon})$ [^29] Here we took $r_{N}=1/\nu=1$. However for the general $r_{N}\neq 0$, we may need to consider $H_{c}=\alpha L_{0}+i\gamma(L_{1}-L_{-1})$ with $d=\alpha^{2}-4\gamma^{2}<0$. This class of Hamiltonian belongs to the ‘heating phase’ in the Floquet CFT language [65],[60]..
We end this section by commenting our take on the questions we have tried to address iff we may think of this deformed CFT quantization as an UV effect:
  • 1.
    What sources the entropy of static patch in the near Nariai reduction to two dimension?
    Thermal entropy of free massless bosons in stretched horizon background. Along the line of [60], one can again think the same as the entanglement entropy of the causal patch vacuum with the outside.  [^30] This is exactly the other side of the quantization contour as described in [60]..
  • 2.
    What is the static patch holography in this context?
    All the entropic degrees of freedom of the extended static patch is restored in the stretched horizon and the Hilbert space is defined by the modular quantization.
  • 2.
    Who is the observer in the story?
    The stretched horizon or the moving membrane defined by the worldsheet deformed CFT on a ring.
6 Discussion
In this note, our main goal is to find possible UV description of effective 2D theory of dilatons arising from reduction of higher dimensional Einstein gravity describing pure and asymptotically dS. While a stable dS2 vacuum is difficult to find in any time-dependent patches of dS2, we have managed to describe a sensible description of (extended and not pure) static patch effective theory in terms of worldsheet CFT with deformed Hamiltonian. Modular quantization of the worldsheet CFT further introduces a stretched horizon with an emergent conformal boundary condition, which acts like an internal observer in this description. It would be definitely interesting to find a set-up where negatively divergent energy density in Milne vacuum could be compensated by gluing it with a very high positive energy density state like large AdS2 blackhole. This could probably lead to some version of cosmology in open universe.
We now discuss some further important points and possible future direction we gained from this note.
[…]
We end our discussion with some speculative remarks on the emergence of worldsheet time. Serious readers may skip this part.
6.1 Comment on emergence of observer time
[…]
$H_{F}=\alpha L_{0}+\beta(L_{1}+L_{-1})+i\gamma(l_{1}-L_{-1})+cc$ (6.1)
[…]
However, the point we wish to emphasize is that, from a UV perspective, the different choices of times (corresponding to different Hamiltonians) should not be a priori preselected, but rather be emergent. In the periodically driven CFT setup[82]-[90], such Hamiltonians appear naturally. For instance, in the simplest scenario one can consider a two-step drive protocol, in which a CFT on a ring is evolved with a Hamiltonian $H_{1}$ for a time interval $T_{1}$, followed by another Hamiltonian $H_{2}$ for a time interval $T_{2}$. This defines an inherently time-dependent system under drive. One can then apply this periodic driving $n$ times, as in:
$\underbrace{e^{-iH_{1}T_{1}}e^{-iH_{2}T_{2}}e^{-iH_{1}T_{1}}e^{-iH_{2}T_{2}}\dots e^{-iH_{1}T_{1}}e^{-iH_{2}T_{2}}}_{\text{n times}}=e^{-iH_{F}t}$ (6.2)
Here $H_{F}$ is the time independent Floquet or effective Hamiltonian and generates a new stroboscopic time $t$. In this driven CFT set up, one can changes different phases dynamically by changing the frequency and amplitude of the drive parameter. By choosing appropriately $H_{1,2}$ one could end up with $H_{F}$ as in (6.1) [54], [55],[56]. Hence different stable configuration of $\alpha,\beta,\gamma$ should correspond to taking different phases. A detailed study of stability of those dynamical phases could be useful to understand those degenerate limits of charged SdS or other near extremal limits in AdS or flat black holes. For instance, from the phase diagram of [56], one can see the heating phases  [^37] This includes $\alpha^{2}<4\gamma^{2}$ as well as $\alpha=\gamma=0$ are separated from distinct non-heating phases by a phase boundary [^38] This includes $\alpha^{2}=4\beta^{2}$$\gamma=0$..
Nevertheless, the main point we want to make here is that the time $t$ inherent to the Floquet Hamiltonian $H_{F}$, is a stroboscopic time $t=n(T_{1}+T_{2})$. In this Floquet sense, the stroboscopic time is emergent. To be more precise, the Floquet observer will measure stroboscopic time in their own clock which is completely different than the clocks for the observers associated to $H_{1,2}$. This is analogous to find a static patch in an accelerating or time dependent (dS) universe. Consequently, to account for the cost of residing in an accelerating universe, static observers are required to observe features like surrounding cosmological (and black hole) horizons. We know that Floquet Hamiltonians in heating phases also exhibit features such as fixed points. We feel, the gravitational observers measure the stroboscopic Floquet time in the worldsheet description, which is inherently emergent. The principle behind the fine tuning of choosing the Hamiltonians $H_{1,2}$, drive parameters and frequencies still remains elusive to us from a gravitational perspective.
Acknowledgement: The author(SD) of this paper strongly believes that any scientific progress is a collective and collaborative process where the authorship is an effort to consolidate all the ideas in a single manuscript. He would like to thank Bobby Ezhuthachan, Rahul Roy and Krishnendu Sengupta for discussion at the early stage of this project. He also wants to thank collectively Dibya Chakraborty, Anirban Dinda, Arnab Kundu, Parthasarathi Majumdar, Debangshu Mukherjee and Koushik Ray, as well as ChatGPT and Perplexity for related discussions and help. He is also grateful to the co-authors of [60]: Bobby Ezhuthachan, Somnath Porey and Baishali Roy for noticing and emphasizing the paper by Solodukhin [42], which provides an important clue to understand the reduction process of this note. He would also like to acknowledge the organisers of the school “Introductory School on Conformal Field Theories” at IACS(18-22 Aug,2025), which helps him to clarify some basic 2D CFT stuff. The research work of SD is supported by a DST INSPIRE Faculty Fellowship.
Appendix A JT gravity and deformed CFT from dimensional reduction of near extremal black holes in near horizon limit
In [1], we claim that the classical limit of JT gravity coupled to CFT with $\mathcal{O}(1)$ central charge is described by a certain deformed CFT in a prescribed limit. Partial justification of this claim was made by the fact that in the semiclassical limit, the backreaction of the conformal fields can be neglected and the dynamics of JT coupled to matter CFT is boiled down to the dynamics of CFT in the UV frame. We recovered the entropy of the coupled system away from extremality, purely from the entropy of deformed CFT in the prescribed classical limit. Here in this appendix, we provide a concrete set-up for this justification where JT coupled to free bosons with $c=1$ arises naturally from S-wave reduction in the near extremal near horizon limit of 4D RN black holes in AdS [91].
[…]
Consider the following magnetically charged AdS4 black holes with the following metric:
$ds^{2}=-(1-\frac{2G_{4}M}{r}+\frac{4\pi Q_{m}^{2}}{r^{2}}+\frac{r^{2}}{L^{2}})dt^{2}+\frac{dr^{2}}{(1-\frac{2G_{4}M}{r}+\frac{4\pi Q_{m}^{2}}{r^{2}}+\frac{r^{2}}{L^{2}})}+r^{2}d\Omega_{2}^{2}$ (A.1)
Here $M,Q_{m}$ are the mass and magnetic charge of the black hole. For the large AdS black hole with $r_{h}>>L$, the extremal charge and mass become [91]
$Q_{ext}^{2}=\frac{3r_{h}^{4}}{4\pi L^{2}},\;M_{ext}=\frac{2r_{h}^{3}}{G_{4}L^{2}}$ (A.2)
Taking the near horizon limit $r\rightarrow r_{h}$ yields the following [91]
$ds^{2}=-\frac{(r-r_{h})^{2}}{L_{2}^{2}}dt^{2}+\frac{L_{2}^{2}}{(r-r_{h})^{2}}dr^{2}+r_{h}^{2}d\Omega_{2}^{2}+\mathcal{O}\left(\frac{r-r_{h}}{r_{h}}\right)$ (A.3)
Here, the AdS2 radius is $L_{2}=\frac{L}{\sqrt{6}}$. This is also true for flat space black holes in the same limit. To obtain the effective gravitational theory in this limit, we begin with 4D Einstein-Maxwell action with negative cosmological constant $\Lambda=-\frac{3}{L^{2}}$, accompanied with a boundary GHY term:
$I=\frac{1}{16\pi G_{4}}\int_{M_{4}}d^{4}X\sqrt{-\hat{g}_{4}}(\hat{R}_{4}-2\Lambda_{4}-4\pi F^{2})+\frac{1}{8\pi G_{4}}\int_{\partial M_{4}}d^{3}X\sqrt{-\hat{\gamma}_{3}}K_{3}$ (A.4)
Unlike the de-Sitter case, here one needs to add the GHY term where $\hat{\gamma}_{3}$ is the induced metric on the three dimensional asymptotic boundary and $K_{3}$ is the trace of extrinsic curvature of the boundary. We also consider the following class of asymtptotically AdS4 metrics with the following form
$ds_{4}^{2}=g_{\mu\nu}dx^{\mu}dx^{\nu}+r_{h}^{2}\phi(x)d\Omega_{2}(\theta,\chi)^{2}$ (A.5)
Further assuming the metric as a solution of Einstein-Maxwell theory, satisfying the equation of motion $F_{\theta\chi}=Q_{m}\sin\theta$, the 4D action reduces to the following 2D action:
$I=\frac{1}{16\pi G_{2}}\int_{M_{2}}d^{2}x\sqrt{-g_{2}}\left(\phi R_{2}-2\Lambda_{4}\phi+\frac{2}{r_{h}^{2}}+\frac{(\nabla\phi)^{2}}{2\phi}-\frac{8\pi Q_{m}^{2}}{r_{h}^{4}\phi}\right)+\frac{1}{8\pi G_{2}}\int_{\partial M_{2}}\sqrt{-\gamma}K$ (A.6)
Here $\frac{1}{G_{2}}\equiv\frac{r_{h}^{2}\Omega_{2}}{G_{4}}$. We note that the total derivative terms from the bulk action gets cancelled from a similar term appearing in the boundary action upon reduction [29]. Again by performing a Weyl rescaling $g_{\mu\nu}\rightarrow\omega^{2}(x)\bar{g}_{\mu\nu}$, we end up with
$I$ $=\frac{1}{16\pi G_{2}}\int_{\bar{M}_{2}}d^{2}x\sqrt{-\bar{g}_{2}}\left(\phi\bar{R}_{2}+\frac{2}{\omega}(\nabla_{\mu}\phi)(\nabla^{\mu}\omega)-2\Lambda\phi\omega^{2}+\frac{2\omega^{2}}{r_{h}^{2}}+\frac{(\nabla\phi)^{2}}{2\phi}-\frac{8\pi Q_{m}^{2}\omega^{2}}{r_{h}^{4}\phi}\right)$
$\;\;\;\;+\frac{1}{8\pi G_{2}}\int_{\partial\bar{M}_{2}}\sqrt{-\bar{\gamma}}\phi_{b}\bar{K}$ (A.7)
Once again a total derivative term from the bulk action $2\nabla^{\mu}\left(\frac{\phi}{\omega}\nabla_{\mu}\omega\right)$ is canceled by a similar term in the boundary action under rescaling. By choosing $\omega\propto\frac{1}{\phi}$ and taking the near horizon limit by expanding $\phi=\phi_{0}+\Phi$ one end up with a standard JT action with boundary and topological Einstein-Hilbert term [91]. However, as in section 2.1, we choose $\omega=c\phi^{-\frac{1}{4}}e^{\beta\phi}$ and obtain the following:
$I$ $=\frac{1}{16\pi G_{2}}\int_{\bar{M}_{2}}d^{2}x\sqrt{-\bar{g}_{2}}\left(\phi\bar{R}_{2}+2\beta(\nabla\phi)^{2}+2\Lambda c^{2}e^{2\beta\phi}\left(\frac{1}{\sqrt{\phi}}-\sqrt{\phi}\right)-\frac{8\pi Q_{m}^{2}}{r_{h}^{4}}c^{2}\phi^{-\frac{3}{2}}e^{2\beta\phi}\right)$
$\;\;\;\;+\frac{1}{8\pi G_{2}}\int_{\partial\bar{M}_{2}}\sqrt{-\bar{\gamma}}\phi_{b}\bar{K}$ (A.8)
Expanding $\phi=\phi_{0}+\Phi$ around $\phi_{0}=1$, we have
$2\Lambda c^{2}e^{2\beta\phi}\left(\frac{1}{\sqrt{\phi}}-\sqrt{\phi}\right)=-2\Lambda c^{2}e^{2\beta}\Phi+\mathcal{O}(\Phi^{2}),\;\text{and}\;c^{2}\phi^{-\frac{3}{2}}e^{2\beta\phi}=c^{2}e^{2\beta}\left[1+(2\beta-\frac{3}{2})\Phi+\mathcal{O}(\Phi^{2})\right].$ (A.9)
Choosing $\beta=\frac{1}{4}$, $c=e^{-\frac{1}{4}}$ and expanding upto linear order in $\Phi$, the action reduces to
$I=$ $\frac{1}{16\pi G_{2}}\int_{\bar{M}_{2}}d^{2}x\sqrt{-\bar{g}_{2}}\left[\frac{1}{2}(\nabla\Phi)^{2}+\Phi\left(\bar{R}_{2}-2\Lambda+\frac{8\pi Q_{m}^{2}}{r_{h}^{4}}\right)\right]+\frac{1}{8\pi G_{2}}\int_{\partial\bar{M}_{2}}\sqrt{-\bar{\gamma}}\Phi_{b}\bar{K}$
$+\frac{\phi_{0}}{16\pi G_{2}}\left(\int_{\bar{M}_{2}}d^{2}x\sqrt{-\bar{g}_{2}}\bar{R}_{2}+2\int_{\partial\bar{M}_{2}}\sqrt{-\bar{\gamma}}\bar{K}\right)-\frac{\phi_{0}Q_{m}^{2}}{2r_{h}^{4}G_{2}}\int_{\bar{M}_{2}}d^{2}x\sqrt{-\bar{g}_{2}}$ (A.10)
$\equiv$ $I_{0}+I_{1}+I_{2}+I_{3}$ (A.11)
Now putting the extremal limit or substituting $Q_{m}$ by $Q_{ext}$, we can see
$-2\Lambda+\frac{8\pi Q_{ext}^{2}}{r_{h}^{4}}=\frac{12}{L^{2}}=\frac{2}{L_{2}^{2}}=-2\Lambda_{2}$ (A.12)
As we argued previously, we have freedom to choose $\bar{g}_{\mu\nu}$ to be locally AdS2 which is Weyl transformed to $g_{\mu\nu}$ with our chosen Weyl factor $\omega$. Since we are interested to study the near extremal and near horizon limit of (A.1), we choose $\bar{g}_{\mu\nu}$ to be in the form of (A.3).
$ds^{2}_{\bar{g}}=-\frac{(r-r_{h})^{2}}{L_{2}^{2}}dt^{2}+\frac{L_{2}^{2}}{(r-r_{h})^{2}}dr^{2}=-4L_{2}^{2}\frac{dz_{+}dz_{-}}{(z_{+}-z_{-})^{2}},\;\text{where}\;z_{\pm}=t\pm L_{2}u,u=\frac{L_{2}}{r-r_{h}}.$ (A.13)
For this metric $\bar{R}_{2}=-\frac{2}{L_{2}^{2}}$. Hence $I_{0}$ reduces to free massless scalar field action in the AdS2 Poincare background. $I_{2}$ is the topological term $I_{top}$ and it provides the extremal entropy of the 4D black hole. $I_{3}$ is the volume term which can be dropped by using suitable counter term. The boundary term $I_{1}$ should also be regulated by using a counter term. Hence the final form reduces to
$I=\frac{1}{16\pi G_{2}}\int_{AdS_{2}}d^{2}z\sqrt{-g_{\{z_{+},z_{-}\}}}\frac{(\nabla_{\{z_{+},z_{-}\}}\Phi)^{2}}{2}+\frac{1}{8\pi G_{2}}\int_{\partial AdS_{2}}\sqrt{-\bar{\gamma}}\Phi_{b}(\bar{K}-1)+I_{top}$ (A.14)
In this rigid AdS2 frame, the boundary term captures the asymptotic symmetries and it’s breaking as described in [92]. To extract the finite term, the usual procedure is to cut-off the rigid AdS near the boundary with a fixed proper length of the boundary term : $g_{\partial AdS}=\frac{1}{\epsilon^{2}}$. The physical boundary time is parametrized by $s$ such that $t=f(s)$ and the fixed proper length boundary condition gives $u=\epsilon f^{\prime}(s)$. However, the solution of the bulk Klein-Gordon equation has both normalizable and non-normalizable modes at the boundary. While the normalizable mode belongs to the solution of physical bulk fields $\Phi$, the non-normalizable mode should be thought of as a source of the boundary term $\Phi_{b}$. In more detail, the free massless scalar field mode solution in AdS2 Poincare is $\Phi(u,t)=e^{i\omega t}\Phi_{\omega}(u)$, where $\Phi_{\omega}=c_{1}\cos[\omega z]+c_{2}\sin[\omega z]$. Hence, at $z\sim 0$, the normalizable solution of $\Phi_{\omega}$ is proportional to $\sin[\omega z]$, whereas the non-normalizable mode is just a constant. Thus without any loss of generality, we can use the following constant boundary condition at the cut-off surface
$\Phi_{u=\epsilon f^{\prime}(s)}=\Phi_{b}=\Phi_{\text{non-normalizable}}=\frac{\Phi_{r}}{\epsilon}$ (A.15)
This particular large value of $\Phi_{b}$ is chosen to extract the finite boundary term as in [92]  [^39] Note that, in the standard JT-dilaton gravity, the solution of the equation of motion incorporates a divergent behavior of the dilaton at the asymptotic boundary. In such a case, the dilaton has no normalizable bulk solution.. Finally using these two boundary conditions, one can extract the boundary action in terms of Schwarzian mode [93]:
$\frac{1}{8\pi G_{2}}\int_{\partial AdS_{2}}\sqrt{-\bar{\gamma}}\Phi_{b}(\bar{K}-1)=\frac{\Phi_{r}}{8\pi G_{2}}\int ds\;Sch\{f(s),s\}$ (A.16)
Since we paramterize the boundary action in terms of boundary or UV time $s$, correspondingly the bulk action should also be modified in the proper boundary time frame. We may consider the following bulk coordinate transformation such that
$z_{\pm}=f(\omega_{\pm}),\;\omega_{\pm}=s\pm\theta\;\text{such that at}\;\theta=\epsilon,\;t=f(s),u=\epsilon f^{\prime}(s).$ (A.17)
Correspondingly the metric transforms as:
$ds^{2}=-4L_{2}^{2}\frac{dz_{+}dz_{-}}{(z_{+}-z_{-})^{2}}\rightarrow-4L_{2}^{2}\left(\frac{\partial z_{+}}{\partial\omega_{+}}\right)\left(\frac{\partial z_{-}}{\partial\omega_{-}}\right)\frac{d\omega_{+}d\omega_{-}}{(z_{+}(\omega_{+})-z_{-}(\omega_{-}))^{2}}$ (A.18)
Since the classical free massless scalar is Weyl invariant, the bulk action is also invariant in the new frame. Hence the total action becomes:
$I=\frac{1}{16\pi G_{2}}\int d^{2}\omega\sqrt{-g_{\{\omega_{+},\omega_{-}\}}}\frac{(\nabla_{\{\omega_{+},\omega_{-}\}}\Phi)^{2}}{2}+\frac{\Phi_{r}}{8\pi G_{2}}\int ds\;Sch\{f(s),s\}+I_{top}$ (A.19)
We may think of this action as Schwarzian coupled to $c=1$ conformal matter. However a crucial difference exists between this and the standard description. In the standard description of JT coupled to conformal matter, the external conformal fields does not couple to dilaton and the classical equation of motion relates the dilaton with the conformal fields. However the dilaton boundary condition reduces the equation of motion to a constraint equation connecting $\frac{d(Sch\{f(s),s\})}{ds}$ to $T_{s\theta}$ at $\theta=0$. Taking $G_{2}\rightarrow 0$ with $cG_{2}<<1$, the constraint gives $Sch\{f(s),s\}$ to be a constant. Hence the on-shell action gives constant Schwarzian piece which is dominated by AdS2 black hole solution.
On the other hand, in the present picture of (A.19), the matter itself plays the role of dilaton and it’s non normalizable mode provides the source of the Schwarzian action. Hence the equation of motion obtained by varying the action with respect to $f(s)$, one again find $Sch\{f(s),s\}$ to be constant. Thus the classical action reduces to the free scalar field in AdS2 black hole background with topological part and constant Schwarzian part. In [1], we have argued the deformed CFT with a specific Hamiltonian generates the UV time $s$ of one-sided AdS2 black hole, for which the thermal entropy [^40] Since the deformed Hamiltonian is identified as the CFT modular Hamiltonian on UHP, the same entropy measures the entanglement entropy of an interval that covers a spatial slice of one-sided AdS2 black hole. contributes to the desired near extremal entropy [^41] Neglecting the extremal entropy part coming from $I_{top}$. in the classical limit. In particular it was found that the entropy $S=\frac{\Phi_{r}}{8\pi G_{2}}Sch\{f(s).s\}=\frac{c\Lambda}{12}Sch\{f(s),s\}$, with the Schwarzian to be constant [^42] Here $c=1$.. Here the stretched horizon cut-off $\Lambda$ plays the role of $\frac{\Phi_{r}}{G_{2}}$ and the semiclassical limit $\Phi_{r}\rightarrow\infty$ or $G_{2}\rightarrow 0$ is directed by taking $\Lambda\rightarrow\infty$. To be more precise, from the effective description of near extremal near horizon physics as free dilatonic CFT, one need to account the factor $\frac{1}{16\pi G_{2}}$ within the deformed Hamiltonian. Hence, more correctly, the spatial IR cut-off $\Lambda$ of deformed CFT plays the role of $\Phi_{r}$ upto some proportionality constant [^43] This differs from the identification of $\Lambda$ with $\frac{1}{G_{2}}$ upto some proportionality constant as proposed in [1].. Hence the classical description of Schwarzian coupled to $c=1$ CFT as described by (A.19), has a direct correspondence to classical limit of the same CFT with a deformed Hamiltonian [1]. To summarize we may think of this in two ways:
  • Either, we may think the worldsheet deformed CFT is dual to the boundary Schwarzian theory with similar density of states(with proper energy rescaling).
  • Or, the JT coupled to free bosonic CFT with Dirichlet boundary condition on scalar fields, is described by the same worldsheet deformed CFT.
In either description, the Hilbert space of modular quantization or the Hilbert space of free scalars in stretched horizon background of AdS2 black hole provides the UV Hilbert space. Perhaps, the 3d version of this description has been realized in [36] [^44] I thank Chethan Krishnan for emphasizing this possibility at some point while discussing [36]..